OpenStudy (anonymous):

what is the slope? (-11.6,19.8) and (11.4)(14.9)

OpenStudy (anonymous):

(a,b) (c,d) slope = (a-c)/(b-d)

OpenStudy (anonymous):

or, that's not right. It's (b-d)/(a-c)

OpenStudy (anonymous):

y=mx+b i believe

OpenStudy (anonymous):

That won't get you the slope, you need to find (the difference in y)(the difference in x) which gives the "rise over run"

OpenStudy (anonymous):

this is alll it says!!!

OpenStudy (anonymous):

so would that be consider not found

OpenStudy (anonymous):

\[m=\frac{x _{1}-x _{2}}{y _{1}-y _{2}}\]

OpenStudy (anonymous):

The ys go on top and xs on bottom

OpenStudy (anonymous):

-23 over 4.9

OpenStudy (anonymous):

I wrote the equation wrong, sorry. the -23 goes in the denomintor and the 4.9 in the numerator

OpenStudy (anonymous):

m=\frac{y _{1}-y _{2}}{x _{1}-x _{2}}

OpenStudy (anonymous):

\[m=\frac{y _{1}-y _{2}}{x _{1}-x _{2}}\]

OpenStudy (anonymous):

-23 on the top and the4.9 on the bottom

OpenStudy (anonymous):

4.9/-23

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

ok